# Symmetry-protection Zeno phase transition in monitored lattice gauge theories

### v1

@@ I Introduction
Lattice gauge theories (LGTs) stand as key models at the crossroads of fundamental physics, condensed matter, and quantum information science [1, 2, 3, 4]. Their unique framework provides a highly rewarding yet challenging target for quantum simulations [5, 6, 7, 8, 9, 10], offering insights into the interplay between highly constrained dynamics [11, 12, 13], possibly governed by nonabelian local symmetries [14, 15], intricate multi-body interactions, and the emergence of topological order with non-Abelian anyons [14, 16, 17, 18, 19, 20]. Quantum simulation protocols, however, are prone to imperfections and approximations that break the gauge symmetries, resulting in errors that lead to nonphysical results. The presence of local conserved quantities in LGTs presents an intriguing avenue for harnessing these symmetries to implement several error-mitigation strategies [21]. For instance, one can modify the Hamiltonian to include energy penalties on unphysical gauge sectors [22, 23, 24, 25, 26, 27, 28, 29] design stochastic processes to destroy phase coherence between physical and nonphysical states [30, 31, 32, 33], or use the gauge symmetry to improve the data quality in post-selection [29, 34]. A particularly appealing avenue is to adapt mid-circuit syndrome measurements, a cornerstone of quantum error-correction and detection algorithms [35], to the context of LTGs, as the local gauge symmetries offer a natural qubit redundancy [36, 37, 38]. So far, however, these proposals assumed fault-tolerant quantum computation or the possibility of performing mid-circuit measurements (together with immediate feedback) arbitrarily many times. However, the effect of the interplay of error strength and the measurement frequency on the system dynamics is still unexplored. A thorough understanding of the resulting stochastic dynamics can, however, prove crucial for developing practical error-mitigation schemes for present-day and near-term quantum simulators, where significant extra resource requirements put full error-correcting capabilities still beyond reach.
@@
In this paper, we show that error protection in digital quantum simulations of LGTs emerges not gradually, but after a critical quantum Zeno transition. This threshold is determined by the ratio between the circuit density of local projective measurements [39, 40, 41] of the gauge symmetry generators and the coherent error strength. Framing our findings within the spectral theory of Liouvillians [42], we highlight the generality of the Zeno transition with respect to different measurement schemes in digital and analog protocols. The presence of this transition has immediate consequences. For example, within the protected Zeno regime, our theory provides a systematic way to perform finite-error scaling to extrapolate exact expectation values. Importantly, the threshold appears already by only passively monitoring the gauge symmetry, which can be done via ancillas also as a delayed measurement, and which therefore is considerably simpler than active error correction. Our results therefore combine insights into stochastic criticality of monitored highly constrained systems with concrete consequences for state-of-the-art quantum simulation protocols.
@@
Our focus lies on the effect of the measurement of local gauge symmetry generators, in close analogy to stabilizer measurements in the framework of error correction [36, 37] or state preparation [65, 66], and the resulting quantum Zeno stochastic dynamics. When these measurements are performed during the time evolution and not only at the end, we effectively realize a dynamical post-selection (DPS) protocol, allowing for both the detection and suppression of gauge violations in digital quantum simulations of LGTs. We illustrate our findings through the paradigmatic example of a $1+1$d ${\mathbb{Z}}_{2}$ LGT with fermionic matter, where we explore the synergistic relationship between stochastic monitored quantum dynamics, error correcting codes [36, 16, 37, 67, 68], and strong constraints.
@@
Moreover, the DPS scheme we implement for digital quantum simulations can be extended to include an on-chip feedback error correction mechanism. This protocol efficiently compensates even for a dissipative error channel, and by using the natural redundancy of the gauge symmetries alone. As we show, the protected phase persists despite the considerably more complex scenario that combes both coherent and incoherent error sources. This finding is significant, as previous protocols protect against coherent errors or sufficiently colored noise, but not white noise [69].
@@
The rest of the paper is organized as follows. In Sec. II, we describe the ${\mathbb{Z}}_{2}$ LGT we study to illustrate the error mitigation schemes. We analyze the gauge protection in discrete-time dynamics in Sec. III and connect it to the quantum Zeno protection in analog quantum simulations. In particular, discuss the mapping emergence of a quantum Zeno transition in Sec. III.1 and analyze the different stochastic unravelings in Sec. III.2. In Sec. IV, we extend the DPS protocol to correct for the effect of incoherent bit-flip noise. We discuss our results and future outlooks in Sec. V. Further details and numerical results are presented in the Appendix.
@@ II Model Hamiltonian
$H_{\rm err}=\lambda_{1}\sum_{n}\tau^{x}_{n,n+1}+\lambda_{2}\sum_{n}\left(\sigma^{+}_{n}\sigma^{-}_{n}+{\rm h.c.}\right)\ ,$ (2)
@@
We consider a digital quantum simulation protocol where the state evolves according to a standard first-order Trotter discretization of the time evolution operator $\textrm{e}^{-iHt}=\left[\textrm{e}^{-iH\Delta t}\right]^{n_{t}}$, with $t=\Delta t\,n_{t}$ being the total evolution time and $n_{t}$ the number of Trotter steps. Each element is further decomposed in the product of simpler unitary operators generated by the single terms in the sums in Eqs. (1) and (2). On top of this, we can consider incoherent noise effects by adding jump operators associated with dissipation or dephasing errors occurring with a certain probability within each time step. We will later consider spin-flip errors in Sec. IV, where we employ a feedback correction mechanism. For now, we focus on the situation where the only error sources are coherent, with a single parameter $\lambda_{1}=\lambda_{2}=\lambda$ setting their strength.
To detect gauge violations arising during the time evolution, we couple the symmetry operators of the model, i.e., the local gauge charges, to ancillary qubits [77, 36], initially prepared in the fiducial state $|0\rangle_{a}$. Mid-circuit measurements on these auxiliary registers project the state either back in the target gauge sector, effectively removing the accumulated errors, or in an orthogonal sector. When a gauge violation is measured, i.e., the ancilla state has flipped, the simulation is discarded as it gives an unphysical contribution to the evaluation of expectation values. The outcome of each measurement is stochastic, despite the discretized evolution being deterministic, with the probability of their outcome depending on the amplitude of the gauge violation. Importantly, we may dilute the measurements, either by measuring only a subset of local charges at each time step or by reducing the frequency of the measurements. This allows for the tuning of the effective protection strength.
@@ III.3 DPS survival probability
where $t=n_{t}\Delta t$ is the evolution time. It decreases faster for larger systems and longer evolution times, besides larger error strength, while it benefits from smaller Trotter steps. If $\Delta t$ is small enough, meaning that the frequency of measurements is faster than the growth rate of the error $\lambda^{2}$, the dynamics reach a quantum Zeno regime where the exponential drift out of the target sector is well approximated by a linear dependence on time $P_{s}(t)\simeq 1-(\lambda^{2}\Delta t)Nt$.
@@
IV Error correction
@@
As we have seen in the previous section, reading out the violations of gauge symmetry at each time step suppresses errors due to a quantum Zeno effect. In addition, it also opens up the possibility to actively correct the measured errors. This active correction scheme is represented by the full circuit sketched in Fig. 1(a). Its implementation requires on-chip feedback during the quantum algorithm. Such a feature is currently actively researched due to its importance for general error correction and is expected to become available in the next generation of quantum chips [81, 82, 83]. We will see below that one can exploit it for significantly improving the quality of gauge theory quantum-simulation data, without the requirement for full error correction capabilities.
@@
First, we test the correction capabilities on a bit-flip (Pauli-X) channel, where we only consider the incoherent errors that the local charge measurements can correct without recurring to more sophisticated QEC protocols [37]. We model this bit-flip error by assuming that, during each Trotter step, an incoherent spin flip can occur on any qubit with probability $p_{\rm err}$. These errors can be detected by local charge measurements and comparing them with the known target sector [36, 77], using the same circuit as sketched in Fig. 1(a). Then, we apply the reverse bit-flip operation and correct the error depending on which gauge symmetries have been violated. This procedure works unambiguously only when the error probability is sufficiently small such that the probability of two errors occurring in the same Trotter step is negligible. Indeed, when two bit-flips affect neighboring local charges, a simple measurement can not determine which qubits were affected by the dissipation errors. In our numerics, when errors with ambiguous detection happen, we discard the run and do not use it to compute the expectation values of observables of interest.
@@
$p_{2}=p_{\rm err}^{2}(1-p_{\rm err})^{L-2}7(N-1)\ ,$ (13)
where $L$ is the total number of qubits (matter and gauge sites), while $N$ is the number of matter sites. The factor $7(N-1)$ is the number of qubit pairs where errors can occur and affect two consecutive local charges. In this estimate, we neglect the effect of the boundaries. Neglecting the possibility that three or more consecutive errors happen in the same Trotter step, the probability that a simulation fails due to these non-correctable errors is, therefore, $1-P_{\rm corr}=1-(1-p_{2})^{n_{t}}$. For the data presented in Fig. 5, where $p_{\rm err}=0.02$, $L=17$, and $n_{t}=20$, this estimate gives a total error probability at the end of the time evolution of $P_{\rm err}=1-P_{\rm corr}\simeq 0.3$. This estimate is compatible with the numerical data: out of 100 trajectories, 16 are discarded because of ambiguous errors (dashed red lines), and 15 lead to wrong results because of nondetectable double errors (outlier grey lines).
@@
$\frac{P_{\rm corr}}{P_{\rm succ}}=\left[\frac{1-p_{\rm err}^{2}(1-p_{\rm err})^{L-2}7(N-1)}{(1-p_{\rm err})^{L}}\right]^{n_{t}}\ .$ (14)
@@
Finally, let us combine the two types of error sources we discussed so far, i.e., local incoherent spin flips and the coherent error driven by the Hamiltonian in Eq. (2). Now, enforcing gauge protection becomes more complicated as bit-flip corrections are (wrongly) applied when either an ancilla measurement or a random spin flip generates a gauge violation.
We adopt the same protocol as discussed so far, with both DPS and bit-flip correction, on top of which we implement a simple post-selection of the corrected trajectories. To eliminate automatically those instances where the correction failed, we exclude the erratic trajectories far from the exact dynamics. First, we evaluate the mean and standard deviation (std) of an observable, the electric Hamiltonian $H_{f}$ for instance, using all the simulations in which the gauge violation has been corrected. Then, we exclude all the data that wander sufficiently far from the average, assuming they are affected by undetectable errors. From the remaining trajectories, we extract a new ensemble average and variance. We iterate this procedure twice, first by excluding data that differ more than one standard deviation from the average, then those that differ more than two updated std from the new mean. We present more details on this technique in App. C. It is important to remark that this post-selection may fail in actual experiments as it requires an accurate estimate of the expectation value of an observable on a single trajectory. Nevertheless, its success suggests that more refined clustering methods might produce similar effects on data collected with projective measurements on the computational basis.
@@
V Conclusions

### v3

@@ I Introduction
Lattice gauge theories (LGTs) stand as key models at the crossroads of fundamental physics, condensed matter, and quantum information science [1, 2, 3, 4]. Their unique framework provides a highly rewarding yet challenging target for quantum simulations [5, 6, 7, 8, 9, 10], offering insights into the interplay between highly constrained dynamics [11, 12, 13], possibly governed by nonabelian local symmetries [14, 15], intricate multi-body interactions, and the emergence of topological order with non-Abelian anyons [14, 16, 17, 18, 19, 20]. Quantum simulation protocols, however, are prone to imperfections and approximations that break the gauge symmetries, resulting in errors that lead to nonphysical results. The local conserved quantities in LGTs present an intriguing avenue for harnessing these symmetries to implement several error-mitigation strategies [21, 22]. For instance, one can modify the Hamiltonian to include energy penalties on unphysical gauge sectors [23, 24, 25, 26, 27, 28, 29, 30] design stochastic processes to destroy phase coherence between physical and nonphysical states [31, 32, 33, 34], or use the gauge symmetry to improve the data quality in post-selection [30, 35, 36].
A particularly appealing approach is to adapt mid-circuit syndrome measurements, a cornerstone of quantum error-correction and detection algorithms [37], to the context of LGTs, as the local gauge symmetries offer a natural qubit redundancy [38, 39, 40, 41]. So far these proposals assumed fault-tolerant quantum computation or the possibility of performing mid-circuit measurements (together with immediate feedback) arbitrarily many times. The effect on the system dynamics of the interplay of error strength and the measurement frequency, also taking into account the further noise they might introduce, is still unexplored. A thorough understanding of the resulting stochastic dynamics can prove crucial for developing practical error-mitigation schemes for present-day and near-term quantum simulators, where significant extra resource requirements put full error-correcting capabilities still beyond reach.
@@
In this paper, we show that error protection in digital quantum simulations of LGTs emerges not gradually, but after a critical quantum Zeno transition. This threshold is determined by the ratio between the circuit density of local projective measurements [42, 43, 44] of the gauge symmetry generators and the coherent error strength. Framing our findings within the spectral theory of Liouvillians [45], we highlight the generality of the Zeno transition with respect to different measurement schemes in digital and analog protocols. The presence of this transition has immediate consequences. For example, within the protected Zeno regime, our theory provides a systematic way to perform finite-error scaling to extrapolate exact expectation values. Importantly, the threshold appears already by only passively monitoring the gauge symmetry, which can be done via ancillas also as a delayed measurement, and which therefore is considerably simpler than active error correction. This measurement-induced protection also holds when taking into account weak depolarizing noise induced by the mid-circuit monitoring, where the Zeno phase survives for a finite range of measurement rates. The combination of mid-circuit measurements with active feedback highlights the potential and limitations of rudimentary quantum error correction in quantum simulation protocols with both unitary and dissipative error channels. Our results combine insights into stochastic criticality of monitored highly constrained systems with concrete consequences for state-of-the-art quantum simulation protocols.
@@
We focus on the effect of the measurement of local gauge symmetry generators, in close analogy to stabilizer measurements in the framework of error correction [38, 39, 41] or state preparation [68, 69], and the resulting quantum Zeno stochastic dynamics. When these measurements are performed during the time evolution and not only at the end, we effectively realize a dynamical post-selection (DPS) protocol, allowing for both the detection and suppression of gauge violations in digital quantum simulations of LGTs. We illustrate our findings mainly through the paradigmatic example of a $1+1$d ${\mathbb{Z}}_{2}$ LGT with fermionic matter, where we explore the synergistic relationship between stochastic monitored quantum dynamics, error correcting codes [38, 16, 39, 70, 71], and strong constraints. We demonstrate the generality of our findings by analysing also gauge theories with $\mathbb{Z}_{3}$ and truncated $U(1)$ symmetry.
@@
Previous works on using measurements to stabilize local symmetries have mostly focused on formal fault-tolerant frameworks [32, 33, 38, 39, 40, 41], often overlooking realistic quench dynamics and how measurements affect it, or on protecting against purely coherent gauge drifts [21, 28, 29, 30, 31, 32], neglecting scenarios with both unitary and dissipative errors. Our scheme, instead, deals with both coherent and incoherent gauge-breaking errors within the same framework and, importantly, using the same circuit elements. It efficiently compensates for a dissipative error channel, thanks to the natural redundancy of the gauge symmetries, while simultaneously enforcing a quantum Zeno protection against a coherent gauge drift. Hence, the protected phase persists despite this considerably more complex scenario that combines both unitary and dissipative error sources. This finding is significant, as previous protocols protect against coherent errors or sufficiently colored noise, but not white noise [72]. Finally, unlike prior studies on engineered dissipation [31, 73], our approach leverages quantum trajectories, enabling a better reconstruction of the target dynamics and a deeper understanding of the role of measurements themselves.
@@
The rest of the paper is organized as follows. In Sec. II, we describe the ${\mathbb{Z}}_{2}$ LGT we study to illustrate the error mitigation schemes. We analyze the gauge protection in discrete-time dynamics in Sec. III and connect it to the quantum Zeno protection in analog quantum simulations. In particular, discuss the mapping emergence of a quantum Zeno transition in Sec. III.1 and analyze the different stochastic unravelings in Sec. III.2. We study the effect of imperfect measurements on the Zeno-protected phase in Sec. IV. In Sec. V, we extend the DPS protocol to correct for the effect of incoherent bit-flip noise. We discuss our results and future outlooks in Sec. VI. Further details and numerical results are presented in the Appendix.
@@ II Model Hamiltonian
$H_{\rm err}=\lambda_{1}\sum_{n}\tau^{x}_{n,n+1}+\lambda_{2}\sum_{n}\left(\sigma^{+}_{n}\sigma^{-}_{n+1}+{\rm h.c.}\right)\ ,$ (2)
@@
We consider a digital quantum simulation protocol where the state evolves according to a standard first-order Trotter discretization of the time evolution operator $\textrm{e}^{-iHt}=\left[\textrm{e}^{-iH\Delta t}\right]^{n_{t}}$, with $t=\Delta t\,n_{t}$ being the total evolution time and $n_{t}$ the number of Trotter steps. Each element is further decomposed in the product of simpler unitary operators generated by the single terms in the sums in Eqs. (1) and (2). On top of this, we can consider incoherent noise effects by adding jump operators associated with dissipation or dephasing errors occurring with a certain probability within each time step. We will later consider spin-flip errors in Sec. V, where we employ a feedback correction mechanism. For now, we focus on the situation where the only error sources are coherent, with a single parameter $\lambda_{1}=\lambda_{2}=\lambda$ setting their strength.
To detect gauge violations arising during the time evolution, we couple the symmetry operators of the model, i.e., the local gauge charges, to ancillary qubits [81, 38, 73], initially prepared in the fiducial state $|0\rangle_{a}$. Mid-circuit measurements on these auxiliary registers project the state either back in the target gauge sector, effectively removing the accumulated errors, or in an orthogonal sector. When a gauge violation is measured, i.e., the ancilla state has flipped, the simulation is discarded as it gives an unphysical contribution to the evaluation of expectation values. The outcome of each measurement is stochastic, despite the discretized evolution being deterministic, with the probability of their outcome depending on the amplitude of the gauge violation. Importantly, we may dilute the measurements, either by measuring only a subset of local charges at each time step or by reducing the frequency of the measurements. This allows for the tuning of the effective protection strength.
@@ III.1 Effective Lindblad master equation and quantum Zeno transition
In Appendix B, we analyze the effective Liouvillian spectrum also for the $\mathbb{Z_{3}}$ and truncated $U(1)$ LGTs, showing that the phenomenology is qualitatively similar and independent from the specific symmetry group.
@@ III.3 DPS survival probability
where $t=n_{t}\Delta t$ is the evolution time. It decreases faster for larger systems and longer evolution times, besides larger error strength, while it benefits from smaller Trotter steps. If $\Delta t$ is small enough, meaning that the frequency of measurements is faster than the growth rate of the error $\lambda^{2}$, the dynamics reach a quantum Zeno regime where the exponential drift out of the target sector is well approximated by a linear dependence on time $P_{s}(t)\simeq 1-(\lambda^{2}\Delta t)Nt$. The decay rate is also connected to the Zeno timescale, corresponding to the simulation time up to which this linear approximation holds and we can extract reliably physical information from gauge-invariant trajectories. This Zeno timescale is approximately $t_{\rm Z}\sim(\lambda^{2}\Delta tN)^{-1}$, meaning it increases linearly with the measurement frequency $\gamma=\frac{1}{2\Delta t}$.
@@
In general, an exponential overhead is common to many error-mitigation techniques [22] and DPS makes no exception. Although this eventually limits their applicability on large systems, direct post-selection methods, for instance, have already been successfully applied to study nontrivial dynamics in LGT experiments [35, 30, 36]. Moreover, the accurate estimate of local observables only requires the system to satisfy the gauge symmetries within their light cone, allowing for extracting meaningful physical information also from trajectories that incur in a small number of gauge violations. Since changing the monitoring frequency allows for tuning the decay rate in Eq. (12), our proposal can prove advantageous both in terms of sampling overhead and protection efficiency with respect to direct-post-selection, making it relevant for system sizes comparable with state-of-the-art NISQ devices ($\sim 100$ qubits), provided the mid-circuit measurements are sufficiently accurate. In the next section, we investigate what happens if this is not the case.
IV noisy measurements
Until now, we have assumed that the gates used to encode the local charge on the auxiliary qubits act with arbitrary precision, without introducing further errors or noise. Despite highlighting the quantum Zeno transition more transparently, this is a significant simplification since two-qubit gates are the main noise sources in realistic hardware.
Figure 5: Magnitude of the real part of Liouvillian eigenvalues as a function of the ratio $\gamma/\lambda$ between the measurement frequency and the coherent error strength, for a measurement fidelity of $\mathcal{F}=0.99$. There remains an intermediate Zeno-protected phase. It is limited from below ($\gamma/\lambda\sim 1$) as before by the build-up of simulation errors when measurements are rare, and from above ($\gamma/\lambda\sim 10$) by large errors accumulating through the measurements themselves once they become too frequent. Marker opacity reflects the density of states and the dashed lines are guides to the eyes to highlight the three competing behaviors.
Here, we relax this assumption and model imperfect measurements through the action of random Pauli noise any time we couple a physical qubit with the ancilla. Specifically, if we measure $G_{n}=-\tau^{z}_{n-1,n}\sigma^{z}_{n}\tau^{z}_{n,n+1}$, we also apply small random rotations around the $\bf{x}$ and $\bf{z}$ axes on the three involved system qubits, given by
$U_{R,n}(\boldsymbol{\xi},\boldsymbol{\eta})=\bigotimes_{k\in\mathrm{supp}(G_{n})}\textrm{e}^{-i\xi_{k}Z_{k}}\textrm{e}^{-i\eta_{k}X_{k}}\ .$ (13)
Figure 6: Relative error of the electric Hamiltonian at times $tJ=10$ (a) and $tJ=20$ (b) for DPS applied with varying measurement intervals $\Delta t_{m}$ and both ideal ($\mathcal{F}=1$) and imperfect ($\mathcal{F}=0.99$) measurements. We compare the data with a post-selection procedure only at the end of the simulation with no intermediate monitoring (horizontal dashed line). At sufficiently large measurement frequencies, DPS improves on top of direct post selection even with noisy measurement circuits. (c) Survival probability as a function of the measurement frequency. A dip appears in correspondence of the Zeno transition, with a worse recovery for finite infidelity $1-\mathcal{F}$. Data for $tJ=20$ and shortest (noisy) measurement period $\Delta t_{m}=\Delta t$ is missing because no physical DPS trajectory was found. Simulation parameters: $N=7$, $\Delta tJ=0.25,f=0.5J$, $\lambda=0.2J$.
$k$ labels the qubits in the support of $G_{n}$ and $Z=\sigma^{z}\ (\tau^{z})$ when $k$ denotes a charge (gauge) degree of freedom, and analogously for $X$. The angles $(\boldsymbol{\xi},\boldsymbol{\eta})$ of the rotations are drawn from a normal distribution with zero mean and a standard deviation $\sigma$ that sets the noise level. $U_{R}$ maintains unitarity of the time evolution on a single trajectory but generates both dephasing and dissipation when we take the ensemble average. We characterize its fidelity through the Hilbert–Schmidt inner product with the identity operator, averaged over noise realizations
$\mathcal{F}=\overline{\left(\mathbb{1},U_{R}(\boldsymbol{\xi},\boldsymbol{\eta})\right)_{\mathrm{HS}}}\simeq 1-3\sigma^{2}\ .$ (14)
The effective error rate induced by noisy measurements is the infidelity $1-\mathcal{F}$.
A first nontrivial question regards the fate of the Zeno transition marked by the branching of the Liouvillian eigenvalues. With finite measurement fidelity, the effective Lindblad equation approximately becomes
$\dot{\rho}$ $=i[\rho,H]+\gamma\sum_{n}\left[G_{n}\rho G^{\dagger}_{n}-\frac{1}{2}\{G^{\dagger}_{n}G_{n},\rho\}\right]$ (15)
$+(1-\mathcal{F})\gamma\sum_{k=1}^{2N-1}\left[L_{k}\rho L^{\dagger}_{k}-\frac{1}{2}\{L^{\dagger}_{k}L_{k},\rho\}\right]\ ,$
where $L_{k}={X_{k},Z_{k}}$ and $k$ runs over all $12N-$ 1 deegrees of freedom. Hence, increasing the measurement frequency also increases the effective error rate. If the eigenvalue branching remains, we expect the presence of an optimal region in the measurement frequency where the Zeno regime still exists. Figure 5, which reports the magnitude of the $\gamma$ dependence of the real part of the Liouvillian eigenvalues for $\lambda=0.2J$ and $\mathcal{F}=0.99$, confirms this picture. The branching that signals the onset of the protected phase at $\gamma/\lambda=1$ is, indeed, still present but the measurement noise prevents the stabilization of the Zeno effect at large frequencies because of a bundle of eigenvalues growing as $-\mathcal{R}(\epsilon_{\mathcal{L}})\propto\gamma$ that “bends” upward the Zeno-protected states. However, a finite range of measurement frequencies remains where the gauge protection from DPS is expected.
To confirm this prediction, we simulate the dynamics of a chain of $N=7$ charge sites ($13$ qubits) with the same quantum quench protocol as described in Sec. III and varying measurement frequency. In practice, we fix the total evolution time $t_{f}J=20$ and the Trotter step $\Delta tJ=0.25$ but we dilute the measurements. We denote with $\Delta t_{m}$ the time interval between consecutive measurement layers, such that the monitoring frequency becomes $\gamma=\frac{1}{2\Delta t_{m}}$. We take $\Delta t_{m}$ as an integer multiple of the Trotter step $\Delta t$.
Figures 6 (a) and (b) report the relative error of the electric field Hamiltonian at two time slices $tJ=10$ (a) and $tJ=20$ (b), averaged over successful DPS trajectories out of 800 independent runs. We compare ideal measurements ($\mathcal{F}=1$, circles) with noisy ones ($\mathcal{F}=0.99$, squares) and a direct post-selection performed only at the end of the run (horizontal dashed line). The latter corresponds to a measurement period of $\Delta t_{m}=tJ$. When the measurements are perfect, increasing their frequency first leads to worse performances with respect to direct post-selection. After the Zeno transition ($2\Delta t_{m}\lambda\simeq 1$), the error starts decreasing and, eventually, DPS performs better than direct post-selection. Interestingly, noisy measurements do not significantly alter this picture and occasionally even lead to better results than ideal DPS. The fluctuations, however, markedly increase for larger measurement frequencies as the accumulated errors tend to diffuse the trajectories around their average while, at the same time, reducing the number of successful ones. At small $\Delta t_{m}$, noise overshadows the Zeno effect and the error starts growing again.
The survival probability also reflects the presence of the Zeno transition and its persistence at a finite error rate in the measurements, as is visible from Fig. 6(c). It shows the survival fraction corresponding to the data in panels (a) and (b) as a function of the measurement frequency. At the Zeno critical point, $P_{s}$ has a clear dip followed by a sharp increase that surpasses the ratio of trajectories retained by direct post-selection (leftmost points). The noisy measurements affect $P_{s}$ in the Zeno regime, reducing the number of trajectories that can be used for reconstructing the target dynamics. In particular, the survival probability in Eq. (12) is modified with an effective error rate that grows with $(1-\mathcal{F})$
$P_{s}(t)\simeq\textrm{e}^{-[\lambda^{2}\Delta t_{m}+(1-\mathcal{F})/\Delta t_{m}]Nt}.$ (16)
Importantly, now it is no longer possible to arbitrarily reduce the exponential decay rate of $P_{s}$ by making more frequent measurements but one has to balance the benefit of the Zeno-like stabilization with the additional noise.
To summarize, imperfect measurements have two main effects: first, they induce an effective relaxation within the target gauge sector, reducing the quality of the reconstructed dynamics. Second, the fraction of successful DPS trajectories decreases due to the higher error rate. Thus, while it is no longer possible to arbitrarily improve the gauge protection by increasing the measurement density, an optimal frequency range exists close to the quantum Zeno transition where mid-circuit measurements still provide the desired mitigation against coherent errors. In this regime, our protocol leads to a better approximation of the target dynamics than standard post-selection, as it suppresses multiple gauge violations that combine in a gauge-invariant error. The existence of the optimal region in the measurement frequency depicted in 6 does not depend on the gauge group. In Appendix B, we report the same phenomenon for the gauge groups $\mathbb{Z}_{3}$ and $U(1)$.
V Error correction
@@
As we have seen in Sec. III, reading out the violations of gauge symmetry at each time step suppresses errors due to a quantum Zeno effect. In addition, it also opens up the possibility to actively correct the measured errors. This active correction scheme is represented by the full circuit sketched in Fig. 1(a). Its implementation requires on-chip feedback during the quantum algorithm. Such a feature is currently actively researched due to its importance for general error correction and is expected to become available in the next generation of quantum chips [85, 86, 87]. We will see below that one can exploit it for significantly improving the quality of gauge theory quantum-simulation data, without the requirement for full error correction capabilities.
@@
First, we test the correction capabilities on a bit-flip (Pauli-X) channel, where we only consider the incoherent errors that the local charge measurements can correct without recurring to more sophisticated QEC protocols [39, 41]. For simplicity, here we resort again to ideal measurements. Alternatively, one could consider a larger error rate incorporating the effect of the two-qubit gates used in the coupling with the ancillas. We model the bit-flip error by assuming that, during each Trotter step, an incoherent spin flip can occur on any qubit with probability $p_{\rm err}$. These errors can be detected by local charge measurements by comparing them with the known target sector [38, 81], using the same circuit as sketched in Fig. 1(a). Then, we apply the reverse bit-flip operation and correct the error depending on which gauge symmetries have been violated. This procedure works unambiguously only when the error probability is sufficiently small such that the probability of two errors occurring in the same Trotter step is negligible. Indeed, when two bit-flips affect neighboring local charges, a simple measurement can not determine which qubits were affected by the dissipation errors. In our numerics, when errors with ambiguous detection happen, we discard the run and do not use it to compute the expectation values of observables of interest.
@@
$p_{2}=p_{\rm err}^{2}(1-p_{\rm err})^{L-2}6(N-1)\ ,$ (17)
where $L$ is the total number of qubits (matter and gauge sites), while $N$ is the number of matter sites. The factor $6(N-1)$ is the number of qubit pairs where errors can occur and affect two consecutive local charges. In this estimate, we neglect the effect of the boundaries. Neglecting the possibility that three or more consecutive errors happen in the same Trotter step, the probability that a simulation fails due to these non-correctable errors is, therefore, $1-P_{\rm corr}=1-(1-p_{2})^{n_{t}}$. For the data presented in Fig. 7, where $p_{\rm err}=0.02$, $L=17$, and $n_{t}=20$, this estimate gives a total error probability at the end of the time evolution of $P_{\rm err}=1-P_{\rm corr}\simeq 0.25$. This estimate is compatible with the numerical data: out of 100 trajectories, 16 are discarded because of ambiguous errors (dashed red lines), and 15 lead to wrong results because of nondetectable double errors (outlier grey lines).
@@
$\frac{P_{\rm corr}}{P_{\rm succ}}=\left[\frac{1-p_{\rm err}^{2}(1-p_{\rm err})^{L-2}6(N-1)}{(1-p_{\rm err})^{L}}\right]^{n_{t}}\ .$ (18)
@@
Finally, let us combine the two types of error sources we discussed so far, i.e., local incoherent spin flips and the coherent error driven by the Hamiltonian in Eq. (2). Now, enforcing gauge protection becomes more complicated as ancilla measurements detect both incoherent spin flips and coherent gauge drift but cannot discriminate between them. A correction with a further bit flip is attempted in both cases but successfully compensates only the former. Correlated coherent errors cannot be inverted, indeed, by a single local bit-flip. The suppression of the coherent gauge drift, therefore, relies on the “passive” protection offered by the quantum Zeno effect through frequent measurements alone.
We adopt the same protocol as discussed so far, with both DPS and bit-flip correction, on top of which we implement a simple post-selection of the corrected trajectories. To eliminate automatically those instances where the correction failed, we exclude the erratic trajectories far from the exact dynamics. First, we evaluate the mean and standard deviation (std) of an observable, the electric Hamiltonian $H_{f}$ for instance, using all the simulations in which the gauge violation has been corrected. Then, we exclude all the data that wander sufficiently far from the average, assuming they are affected by undetectable errors. From the remaining trajectories, we extract a new ensemble average and variance. We iterate this procedure twice, first by excluding data that differ more than one standard deviation from the average, then those that differ more than two updated std from the new mean. We present more details on this technique in App. D. It is important to remark that this post-processing may fail in actual experiments as it requires an accurate estimate of the expectation value of an observable on a single trajectory. Nevertheless, its success suggests that more refined clustering methods might produce similar effects on data collected with projective measurements on the computational basis.
@@
A key take-home message is that within the framework of mid-circuit measurements, which naturally occur in any error-correction scheme, both coherent and incoherent errors can be handled. The efficiency of such protocols is limited by the measurement accuracy and the consequent maximum monitoring frequency. Understanding the intricate interplay between these many ingredients is a crucial step to bridge current NISQ devices to the primitive fault-tolerant ones, where the measurement-induced effects and imperfect error detection could still affect the circuit dynamics in nontrivial ways.
VI Conclusions